OpenStudy (anonymous):

y''-5y' +6y = x how to find particular solution of this dif. equation ? please give detail :) i didnt understand anything :)

OpenStudy (anonymous):

I think there is something wrong with the question as appears to just have one half of the differential equation and has " and ' - not sure if that should be something else. y-5y+6y=x would mean y-11y=x = -10y=x but as you have nothing it equals to it doesn't really help

OpenStudy (anonymous):

Na dude, it's a differential equation, means you have to find such a function y=y(x) which satisfies it. y' and y'' are its derivatives.

OpenStudy (anonymous):

according to middle east technical unuversity 's book particular solution is y = x/6 +5/36 :S

OpenStudy (anonymous):

but i didnt understand how he find this solution

OpenStudy (anonymous):

ah ok - http://www.wolframalpha.com/input/?i=y%27%27-5y%27+%2B6y+%3D+x - wolframalpha pretty useful for helping with equations

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Y(s)=1/(s^2-5s+6)*s^2

OpenStudy (anonymous):

factor it as much as you can and do partial fractions on it, the inverse laplace transform

OpenStudy (anonymous):

then****

OpenStudy (anonymous):

part of your solution will be the solution to the complimentary equation the other part will be the solution to the response of the system due to the forcing function